OpenStudy (anonymous):

a body is projected horizontally with speed 20 m/s from top of a tower . what will be it speed nearly after 5 sec ? g= 10

OpenStudy (compphysgeek):

the body will only experience vertical accelation due to gravity. the horizontal speed therefore remains constant at 20 m/s. vertically the body is released from rest. v = g*t. entered your values into the equation you'll get v after 5 seconds. now you have a body travelling with a horizontal speed of 20 m/s and a vertical speed v (= 50 m/s). Using Pythagoras' Theorem you can calculate the total speed of the body \[ v_{horiz}^2 + v_{vert}^2 = v_{total}^2\]